# Robustness and classical proxy of entanglement in variants of quantum walk

### v1

@@ Abstract
Quantum walk (QW) utilizes its internal quantum states to decide the displacement, thereby introducing single-particle entanglement between the internal and positional degrees of freedom. By simulating three variants of QW with the conventional, symmetric, and split-step translation operators with or without classical randomness in the coin operator, we show the entanglement is robust against both time- and spatially- dependent randomness, which can cause localization transitions of QW. We propose a classical quantity call overlap, which literally measures the overlap between the probability distributions of the internal states, as a proxy of entanglement. The overlap captures the inverse behavior of the entanglement entropy in most cases, which can be explained by analyzing the structure of the total wave function of the walker. We test the limitation of the classical proxy by constructing a special case with high population imbalance between the internal states to blind the overlap. Possible implications and experimental measurements are also discussed.
@@ I Introduction
Direct measurements of entanglement of quantum systems experimentally, however, have proven to be a difficult task. Some forms of proxy quantities have been used to infer the entanglement, such as statistical methods involving charge fluctuation measurements of non-interacting fermions [29], using many-body interference of quantum twins [30, 31], measuring the Renyi entropy [32, 33], and a potential method using machine learning to identify entanglement within quantum states [34]. These methods typically only work for many-body quantum systems and are not well suited or easily applied to systems with SPE. Moreover, they can be experimentally difficult to implement or computationally expensive for larger systems.
Here, we introduce a classical quantity called overlap, which is literally the overlap between the equal-time probability distributions of the internal states of the walker. We will show that the overlap takes on an inverse relationship with the entanglement entropy of the walker, thereby serving as a classical proxy of the entanglement. We also use an analogy to a composite system of two spin-$1/2$ particles to explain where and how entanglement reaches maximum or minimum and why the overlap catches the features. Importantly, the classical nature of the overlap makes it a more experimentally accessible quantity that can be used to investigate the entanglement present in QW. We also present a systematic study of the steady-state entanglement entropy for a variety of QWs with or without classical randomness of the time- or spatially- dependent type. While the overlap works well for most cases, we deliberately study a special case with high population imbalance of the internal states to show when the overlap may not faithfully track the entanglement.
@@ II.2 Symmetric walk
The symmetric walk creates probability distributions that resemble two delta functions moving out evenly, as can be seen in Fig. 2. This two-peak structure is ubiquitous over different $\theta$ values with only the location of the peaks being changed. The lack of spreading and asymmetry of the probability distributions is a signature of the symmetric QW, which will result in low entanglement in our later discussion.
@@ III.1 Time-dependent random coins
The localization of the conventional QW with time-dependent randomness can be seen in Fig. 4 (a), where the probability distributions take the form of Gaussian wave packets when $\Delta\theta$ exceeds the critical value. The distributions of the two internal states, however, do not fully overlap due to the asymmetric translation operators. Meanwhile, the split-step QW in general shows similar behavior as the conventional QW, but the symmetric QW does not localize and maintains the double delta-peak distributions in the presence of time-dependent randomness.
@@ III.2 Spatially dependent random coins
Figure 4: Probability distributions $P_{\pm}$ of the conventional QW with (a) time-dependent, and (b) spatially-dependent randomness in the coin operator. Here $N=1000$, $\theta_{0}=\frac{\pi}{6}$, $\Delta\theta=0.25$ in (a), and $\Delta\theta=0.3$ in (b).
The localization of the conventional QW with spatially-dependent randomness can be seen in Fig. 4 (b), where the probability distributions take the form of localized exponential wave packets when $\Delta\theta$ exceeds the critical value. Again, the distributions of the two internal states do not fully overlap due to the asymmetric translation operators. The split-step QW in general shows similar behavior as the conventional QW. In contrast, the symmetric QW still exhibits its double delta-peak distributions in the presence of spatially dependent randomness and does not localize.
@@ IV Entanglement entropy and its classical proxy
where $\lambda_{\pm}(t)$ are the eigenvalues of the reduced density matrix $\rho_{s}$ for a given step in time. We remark that we follow a logarithm of base $e$, which differ by a factor of $\approx 0.69$ in the values when compared to other works [43, 22, 45] in base $2$. The typical time evolution of the entanglement entropy of QW can be seen in the upper inset of Fig. 5, where the values increases initially but saturates into a steady state. We then evaluate the steady-state value by averaging the long-time behavior to ensure the result is free from transient effects.
@@ IV.1 Overlap as classical proxy of entanglement
$\mathcal{O}(x,t)=\sum_{x=-N}^{N}P_{+}(x,t)P_{-}(x,t).$ (6)
@@
In other words, the above expression measures the overlap of the probability distributions of the two internal states. The reason the overlap reflects the entanglement is as follows. If we write the time-evolved wavefunction of the walker in the form $|\psi(t)\rangle=|\psi_{+}(x,t)\rangle|+\rangle+|\psi_{-}(x,t)\rangle|-\rangle$, where $|\psi_{\pm}(x,t)\rangle$ represent the wave packets of the $\pm$ components, it resembles a state of two spin $1/2$ systems of the form $\frac{1}{\sqrt{2}}(|s_{a}\rangle_{1}|s_{b}\rangle_{2}+|s_{c}\rangle_{1}|s_{d}\rangle_{2})$. Here the two spins are labeled as $1,2$ and $s_{j}=\pm 1/2$ for $j=a,b,c,d$. The entanglement entropy will reach maximum for the Bell states and minimum for product states [14]. From the analogy, the entanglement of QW will be minimum if the two wave packets $|\psi_{\pm}(x,t)\rangle$ have maximal overlap in real space ($|\psi_{\pm}\rangle\approx|\bar{\psi}\rangle$) because the total wavefunction will have the product-state form $|\bar{\psi}\rangle(|+\rangle+|-\rangle)$. In contrast, if the wave packets of the internal states are well separated in real space with minimal overlap, we may treat $|\psi_{+}\rangle$ as |"+"ket""|"+$\rangle$| " + " ⟩ and $|\psi_{-}\rangle$ as |""ket""|"-$\rangle$| " - " ⟩ in real space, so the total wave function has the Bell-state form (|"+"|++|""|)ket""ketket""ket(|"+$\rangle$|$\rangle$+|"-$\rangle$|$\rangle$)( | " + " ⟩ | + ⟩ + | " - " ⟩ | - ⟩ ) with maximal entanglement. Therefore, the overlap indeed serves as an inverse indicator of the entanglement of QW. Our numerical results will confirm the validity of the relation except in some special cases.
@@
The typical time evolution of the overlap of QW is shown in the lower inset of Fig. 5. We note that while the overlap does provide an accessible experimental proxy to indicate when a given quantum walk has finite entanglement entropy, it could require a longer time before the overlap settles into a steady state. Moreover, the overlap may have some large spikes within the first few steps of the walk but eventually decays into its steady-state value at longer times. We will show the steady-state value of the overlap by taking the average at long times to ensure the result is not influenced by transient values.
@@ V.1 QW without classical randomness
Figure 5: Steady-state entanglement entropy $\overline{ES}$ (top panel) and overlap $\mathcal{O}$ (bottom panel) for the three quantum walks without classical randomness. Here $N=500$, and CW, SSW, and SW refer to the conventional walk, split-step Walk, and symmetric walk denoted by the dashed, dotted, and dot-dash lines, respectively.
@@ V.2 Time-dependent random coins
We now introduce classical randomness via time-dependent random rotation operators to the three types of QW. In Figure 6, we plot the steady-state values of the entanglement entropy $(\overline{ES})$ and the overlap $(\mathcal{O})$ as functions of the rotation operator angle $\theta$ with fixed $\Delta\theta$. For the QWs with classical randomness, however, we focus the systems in the localization regime by choosing a relatively large $\Delta\theta$. Given the rotation angles are randomly chosen from $\theta\pm\Delta\theta$, we only show the range from $\theta=\Delta\theta$ to $\theta=\pi-\Delta\theta$. Further, because we are interested in the entanglement when the walk is fully localized, we discard the results from the split-step QW near $\theta=0$ and $\pi$ where the walk was not fully localized even with the relative large value of $\Delta\theta$.
@@
Figure 6: Steady-state entanglement entropy $\overline{ES}$ and overlap $\mathcal{O}$ for the three QWs with time-dependent randomness. Here $N=500$ and $\Delta\theta=0.3$. The shaded regions truncate where the split-step walk does not show localization. CW, SSW, and SW refer to the conventional, split-step, and symmetric Walks, respectively.
@@ V.3 Spatially dependent random coins
Figure 7: Steady-state entanglement entropy $\overline{ES}$ and overlap $\mathcal{O}$ for the three QWs with spatially-dependent randomness. Here $N=500$ and $\Delta\theta=0.3$. CW, SSW, and SW refer to the conventional walk, split-step walk, and symmetric Walk, respectively.
@@ VI.1 More about population imbalance
Figure 10: Steady-state entanglement entropy $\overline{ES}$ and overlap $\mathcal{O}$ for the three QWs with $\phi_{1}=\phi_{2}=0$, $N=500$, and no classical randomness. Here CW, SSW, and SW refer to the conventional, split-step, and symmetric QWs, respectively.
The symmetric QW shown in Fig. 10, however, exhibits interesting behavior. One can see that near $\theta=\pi/4,3\pi/4$, the overlap reaches minimum despite vanishing entanglement entropy. Therefore, the overlap does not reflect the lack of entanglement at those particular values of $\theta$. The reason for the failure of the overlap to indicate the entanglement around those special values can be understood by examining the probability distributions shown in Fig. 9(c). Without the constraint of population balance, the symmetric walk can have time instances with extremely high population imbalance, where one internal state has nearly zero population and causes the overlap to decrease substantially according to Eq. (6). Meanwhile, the wave packets of the two internal states still coincide in real space even in the presence of population imbalance, thereby making the total state more like a product state with low entanglement.
Therefore, we found a special case of the symmetric QW around particular values of $\theta$ with high population imbalance due to the choice of $\phi_{1,2}$, where the overlap no longer serves as a faithful proxy of the entanglement entropy. This observation demonstrates the limitation of classical proxies, such as the overlap, of genuine quantum objects, such as the entanglement entropy. Nevertheless, the overlap has been shown to work in the rest of the cases, so one only needs to exercise caution when high population imbalance of the internal states is encountered.
@@ VI.2 Implications
[…]

### v2

@@ Abstract
Quantum walk (QW) utilizes its internal quantum states to decide the displacement, thereby introducing single-particle entanglement between the internal and positional degrees of freedom. By simulating three variants of QW with the conventional, symmetric, and split-step translation operators with or without classical randomness in the coin operator, we show the entanglement is robust against both time- and spatially- dependent randomness, which can cause localization transitions of QW. We propose a classical quantity call overlap, which literally measures the overlap between the probability distributions of the internal states, as a proxy of entanglement. The overlap is associated with the off-diagonal terms of the reduced density matrix in the internal space, which then reflects its purity. Therefore, the overlap captures the inverse behavior of the entanglement entropy in most cases. We test the limitation of the classical proxy by constructing a special case with high population imbalance between the internal states to blind the overlap. Possible implications and experimental measurements are also discussed.
@@ I Introduction
Direct measurements of entanglement of quantum systems experimentally, however, have proven to be a difficult task. Some forms of proxy quantities have been used to infer the entanglement, such as statistical methods involving charge fluctuation measurements of non-interacting fermions [29], using many-body interference of quantum twins [30, 31], measuring the Renyi entropy [32, 33], extracting the eigen-spectrum of the correlation matrix [34], and a potential method using machine learning to identify entanglement within quantum states [35]. These methods typically only work for many-body quantum systems and are not well suited or easily applied to systems with SPE. Moreover, they can be experimentally difficult to implement or computationally expensive for larger systems. Recently, there has been progress in transmitting entangled states through quantum walks with non-Hermitian or topological properties [36, 37, 38]. These results show further applications of quantum walks.
Here, we introduce a classical quantity called overlap, which is literally the overlap between the equal-time probability distributions of the internal states of the walker. We will show that the overlap takes on an inverse relationship with the entanglement entropy of the walker, thereby serving as a classical proxy of the entanglement. We use an explicit construction of the total and reduced density matrices to show that the off-diagonal terms are associated with the overlap, which in turn increases with the purity of the reduced density matrix and determines the entanglement of the total wavefunction. Moreover, we also present an analogy to a composite system of two spin-$1/2$ particles to further explain where and how the entanglement reaches maximum or minimum and why the overlap catches the features. Importantly, the classical nature of the overlap makes it a more experimentally accessible quantity that can be used to investigate the entanglement present in QW. We also present a systematic study of the steady-state entanglement entropy and overlap for a variety of QWs with or without classical randomness of the time- or spatially- dependent type. While the overlap works well for most cases, we deliberately study a special case with high population imbalance of the internal states to show when the overlap may not faithfully track the entanglement.
@@ II.2 Symmetric walk
[…]
@@ III.1 Time-dependent random coins
[…]
@@ III.2 Spatially dependent random coins
Figure 4: Probability distributions $P_{\pm}$ of the conventional QW with (a) time-dependent, and (b) spatially-dependent randomness in the coin operator. The two internal states still avoid each other in real space despite the localization. Here $N=1000$, $\theta_{0}=\frac{\pi}{6}$, $\Delta\theta=0.25$ in (a), and $\Delta\theta=0.3$ in (b).
[…]
@@ IV Entanglement entropy and its classical proxy
where $\lambda_{\pm}(t)$ are the eigenvalues of the reduced density matrix $\rho_{s}$ for a given step in time. We remark that we follow a logarithm of base $e$, which differ by a factor of $\approx 0.69$ in the values when compared to other works [47, 22, 49] in base $2$. The typical time evolution of the entanglement entropy of QW can be seen in the upper inset of Fig. 5, where the values increases initially but saturates into a steady state. We then evaluate the steady-state value ($\overline{ES}$) by averaging the long-time behavior to ensure the result is free from transient effects.
@@ IV.1 Overlap as classical proxy of entanglement
$\mathcal{O}(t)=\sum_{x=-N}^{N}P_{+}(x,t)P_{-}(x,t).$ (6)
In other words, the above expression measures the overlap of the probability distributions of the two internal states. The reason the overlap reflects the entanglement is as follows. The time-evolved wavefunction of the walker has the form $|\psi(t)\rangle=|\psi_{+}(x,t)\rangle|+\rangle+|\psi_{-}(x,t)\rangle|-\rangle$, where $|\psi_{\pm}(x,t)\rangle$ represent the wave packets of the $\pm$ components. The total density matrix is then $\rho_{tot}=|\psi(t)\rangle\langle\psi(t)|$. If we use a real-space orthonormal basis $\{|x\rangle\}$, the reduced density matrix in the internal space is $\rho_{s}=Tr_{x}(\rho_{tot})=\sum_{x}\langle x|\rho_{tot}|x\rangle$. By using the resolution of the identity, $I=\sum_{x}|x\rangle\langle x|$, the reduced density matrix becomes
$\rho_{s}$ $=$ $\langle\psi_{+}|\psi_{+}\rangle|+\rangle\langle+|+\langle\psi_{-}|\psi_{+}\rangle|-\rangle\langle+|+$ (7)
$\langle\psi_{+}|\psi_{-}\rangle|+\rangle\langle-|+\langle\psi_{-}|\psi_{-}\rangle|-\rangle\langle-|.$
If $|\psi(t)\rangle$ is entangled (separable), $\rho_{s}$ is a mixed (pure) state with purity $Tr(\rho_{s}^{2})<1$ ($Tr(\rho_{s}^{2})=1$[14]. The entanglement entropy (5) checks the purity of $\rho_{s}$ via the von-Neumann entropy formula and reflects the entanglement of $\rho_{tot}$. However, a direct calculation shows that the purity of $\rho_{s}$ is $Tr(\rho_{s}^{2})=|\langle\psi_{+}|\psi_{+}\rangle|^{2}+|\langle\psi_{-}|\psi_{-}\rangle|^{2}+2|\langle\psi_{+}|\psi_{-}\rangle|^{2}$, which increases with the overlap $\mathcal{O}(t)\approx|\langle\psi_{+}|\psi_{-}\rangle|^{2}$. In this perspective, the physical meaning of the overlap (6) is associated with the magnitude of the off-diagonal terms of $\rho_{s}$, which reflects the purity of $\rho_{s}$ in a similar fashion as the entanglement entropy.
Therefore, the maximal (minimal) overlap implies high (low) purity of $\rho_{s}$, which in turn indicates low (high) entanglement in $|\psi(x,t)\rangle$. For partial overlaps between the two internal states in real space, there is finite entanglement since $\rho_{s}$ is not completely pure in general. The above argument explains why the overlap serves as a classical proxy to reflect the inverse behavior of the entanglement. Our simulations of several variants of QW with or without classical randomness will confirm the usefulness of the overlap as a classical proxy for the entanglement entropy in the steady state. As will be shown shortly, the corresponding entanglement entropy is high. However, we caution that the overlap (6) only involves the measurable densities of the internal states, so there exist situations where it may not faithfully reflect the change in entanglement. We will present a specially designed case to illustrate this subtlety later.
@@
One may also understand the relation between the overlap and entanglement by the resemblance between the state $|\psi(t)\rangle=|\psi_{+}(x,t)\rangle|+\rangle+|\psi_{-}(x,t)\rangle|-\rangle$ and a composite state of two spin-$1/2$ systems of the form $\frac{1}{\sqrt{2}}(|s_{a}\rangle_{1}|s_{b}\rangle_{2}+|s_{c}\rangle_{1}|s_{d}\rangle_{2})$. Here the two spins are labeled as $1,2$ and $s_{j}=\pm 1/2$ for $j=a,b,c,d$. The entanglement entropy of the two-spin state will reach maximum for the Bell states and minimum for product states [14]. From the analogy between the QW state and two-spin state, the entanglement of QW will be minimum if the two wave packets $|\psi_{\pm}(x,t)\rangle$ have maximal overlap in real space ($|\psi_{\pm}\rangle\approx|\bar{\psi}\rangle$) because the total wavefunction will have the product-state form $|\bar{\psi}\rangle(|+\rangle+|-\rangle)$. In contrast, if the wave packets of the internal states are well separated in real space with minimal overlap, we may treat $|\psi_{+}\rangle$ as |"+"ket""|"+$\rangle$| " + " ⟩ and $|\psi_{-}\rangle$ as |""ket""|"-$\rangle$| " - " ⟩ in real space, so the total wave function has the Bell-state form (|"+"|++|""|)ket""ketket""ket(|"+$\rangle$|$\rangle$+|"-$\rangle$|$\rangle$)( | " + " ⟩ | + ⟩ + | " - " ⟩ | - ⟩ ) with maximal entanglement. Therefore, the overlap indeed serves as an inverse indicator of the entanglement of QW. Our numerical results will confirm the validity of the relation except in some special cases.
@@
The typical time evolution of the overlap of QW is shown in the lower inset of Fig. 5. We note that while the overlap does provide an accessible experimental proxy to indicate when a given quantum walk has finite entanglement entropy, it could require a longer time before the overlap settles into a steady state. Moreover, the overlap may have some large spikes within the first few steps of the walk but eventually decays into its steady-state value at longer times. We will show the steady-state value of the overlap $(\overline{\mathcal{O}})$ by taking the average at long times to ensure the result is not influenced by transient values.
@@ V.1 QW without classical randomness
Figure 5: Steady-state entanglement entropy $\overline{ES}$ (top panel) and overlap $\overline{\mathcal{O}}$ (bottom panel) for the three quantum walks without classical randomness. Here $N=500$, and CW, SSW, and SW refer to the conventional walk, split-step walk, and symmetric walk denoted by the dashed, dotted, and dot-dash lines, respectively. The upper (lower) inset shows the behavior of $ES$ ($\mathcal{O}$) as a function of time for the conventional walk with $N=500,\theta=\frac{\pi}{6}$.
@@ V.2 Time-dependent random coins
We now introduce classical randomness via time-dependent random rotation operators to the three types of QW. In Figure 6, we plot the steady-state values of the entanglement entropy $(\overline{ES})$ and the overlap $(\mathcal{\overline{O}})$ as functions of the rotation operator angle $\theta$ with fixed $\Delta\theta$. For the QWs with classical randomness, however, we focus the systems in the localization regime by choosing a relatively large $\Delta\theta$. Given the rotation angles are randomly chosen from $\theta\pm\Delta\theta$, we only show the range from $\theta=\Delta\theta$ to $\theta=\pi-\Delta\theta$. Further, because we are interested in the entanglement when the walk is fully localized, we discard the results from the split-step QW near $\theta=0$ and $\pi$ where the walk was not fully localized even with the relative large value of $\Delta\theta$.
@@
Figure 6: Steady-state entanglement entropy $\overline{ES}$ and overlap $\overline{\mathcal{O}}$ for the three QWs with time-dependent randomness. Here $N=500$ and $\Delta\theta=0.3$. The shaded regions truncate where the split-step walk does not show localization. CW, SSW, and SW refer to the conventional, split-step, and symmetric walks, respectively.
@@ V.3 Spatially dependent random coins
Figure 7: Steady-state entanglement entropy $\overline{ES}$ and overlap $\overline{\mathcal{O}}$ for the three QWs with spatially-dependent randomness. Here $N=500$ and $\Delta\theta=0.3$. CW, SSW, and SW refer to the conventional walk, split-step walk, and symmetric walk, respectively.
@@ VI.1 More about population imbalance
Figure 10: Steady-state entanglement entropy $\overline{ES}$ and overlap $\overline{\mathcal{O}}$ for the three QWs with $\phi_{1}=\phi_{2}=0$, $N=500$, and no classical randomness. Here CW, SSW, and SW refer to the conventional, split-step, and symmetric QWs, respectively.
Figure 11: Steady-state entanglement entropy $\overline{ES}$ and overlap $\overline{\mathcal{O}}$ for the three QWs with time-dependent randomness and $\phi_{1}=\phi_{2}=0$. Like in Figure 6, the shaded regions truncate where the split-step walk does not show localization. CW, SSW, and SW refer to the conventional, split-step, and symmetric walks, respectively. Here $N=500$ and $\Delta\theta=0.3$.
Figure 12: Steady-state entanglement entropy $\overline{ES}$ and overlap $\overline{\mathcal{O}}$ for the three QWs with spatially-dependent randomness and $\phi_{1}=\phi_{2}=0$. Here $N=500$ and $\Delta\theta=0.3$. CW, SSW, and SW refer to the conventional, split-step, and symmetric walks, respectively.
The symmetric QW shown in Fig. 10, however, exhibits interesting behavior. One can see that near $\theta=\pi/4,3\pi/4$, the overlap reaches minimum despite vanishing entanglement entropy. Therefore, the overlap does not reflect the lack of entanglement at those particular values of $\theta$. The reason for the failure of the overlap to indicate the entanglement around those special values can be understood by examining the probability distributions shown in Fig. 9(c). Without the constraint of population balance, the coin operator of the symmetric walk no longer spreads the weights of the internal states evenly when the initial condition is symmetric. Thus, there can be time instances with extremely high population imbalance for some values of $\theta$, where one internal state has nearly zero population. As a consequence, the overlap decreases substantially according to Eq. (6) in such a situation. Meanwhile, the wave packets of the two internal states still coincide in real space due to the equal shifts by the translation operator of the symmetric walk even in the presence of population imbalance, thereby making the total state more like a product state with low entanglement.
Therefore, we found a special case of the symmetric QW around particular values of $\theta$ with high population imbalance due to the choice of $\phi_{1,2}$, where the overlap no longer serves as a faithful proxy of the entanglement entropy. This observation demonstrates the limitation of classical proxies, such as the overlap, of genuine quantum objects like the entanglement entropy. Nevertheless, the overlap has been shown to work in the other cases, so one only needs to exercise caution when high population imbalance of the internal states is encountered.
To complete the discussion on QW with population imbalance between the internal states, we show the steady-state entanglement entropy and overlap for the three variants of QW with $\phi_{1}=0=\phi_{2}$ and time- or spatially- dependent randomness in the coin operator. Figs. 11 and 12 show $\overline{ES}$ and $\overline{\mathcal{O}}$ in the presence of time-and spatially- dependent randomness with population imbalance, respectively. For both the conventional and the split-step walks, we see a behavior that resembles that which was found in the equal-population cases in Figs. 5 and 7 except the fluctuations due to randomness.
In contrast, while the entanglement entropy of the symmetric walk resembles that of the equal-population case even in the presence of time- or spatially-dependent randomness, the overlap exhibits fluctuations due to the randomness. We caution that the two-peak probability distribution of the symmetric walk still remains. However, the weights of the two internal states are constantly varying due to the random coin operator when the equal-population condition no longer holds. On the other hand, the noisy behavior renders the highly population-imbalanced configurations negligible in the symmetric walk with time- or spatially-dependent randomness, thereby restoring the overlap as a classical proxy for the entanglement entropy, as shown in Figs. 11 and 12, even in the presence of population imbalance between the internal states.
@@ VI.2 Implications
Our analysis of the three variants of QW has showed interesting entanglement between the internal and positional degrees of freedom while the classical proxy by the overlap captures the change of entanglement. Experimental realizations of quantum walks have shown measurements of the probability distributions $P_{\pm}$. For example, Ref. [55] implemented the discrete time quantum walk in a photonic system where the horizontal and vertical photon polarization ($|H\rangle,|V\rangle$) corresponds to the internal |+,|ketket|$\rangle$,|$\rangle$| + ⟩ , | - ⟩ states, respectively. The probability amplitudes of different polarization can be measured separately, allowing for a reconstruction of the $P_{\pm}$ distributions. There have also been works investigating optical techniques to preform state-dependent imaging of ultracold atoms [56, 57] and map out the densities of individual internal states. Once the probability distributions $P_{\pm}$ of the internal states are obtained, the overlap can be evaluated by Eq. (6). Moreover, experiments with atomic Bose-Einstein condensates [58] have been used to create both quantum walk and classical randomness introduced through the phase factors of the rotation operator. Additionally, Ref. [59] studied QW using a photonic setup where the rotation operator is applied using purely passive linear optics while Ref. [60] constructed QW on solid-state photonic devices.